Hey, Im stuck on this one question for HOMEWORK, so I would like if this would be answered ASAP.
the question is: The length of a rectangle is ten less than three times its width. If the width is increased by five and
the length is increased by two, then the perimeter of the new rectangle is 90cm. Find the dimensions of
the original rectangle.

Found 2 solutions by solver91311, nerdybill:
Answer by solver91311(24713)


Let represent the width of the original rectangle. Then must represent the length of the original rectangle, must represent the width of the larger rectangle and must represent the length of the larger rectangle.

The perimeter of a rectangle is given by two times the length plus two times the width. Since the perimeter of the new rectangle is 90 we can say:



Solve for , then calculate

John

My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism


Answer by nerdybill(7384)

The length of a rectangle is ten less than three times its width. If the width is increased by five and the length is increased by two, then the perimeter of the new rectangle is 90cm. Find the dimensions of the original rectangle.
.
Let x = width
then
3x-10 = length
.
width increased by 5:
x+5
length increased by 2:
3x-10+2 = 3x-8
.
2(x+5 + 3x-8) = 90
2(4x+5-8) = 90
2(4x-3) = 90
4x-3 = 45
4x = 48
x = 14 cm (width)
.
Length:
3x-10 = 3(14)-10 = 42-10 = 32 cm
.
Dimensions:
14 cm by 32 cm